OpenStudy (anonymous):

log (x-3)= 1- logx how to solve with steps please

OpenStudy (anonymous):

\[\log(x-3)=1-\log(x)\] solve for \(x\) right?

OpenStudy (anonymous):

you might want to start with \[\log(x-3)+\log(x)=1\]

OpenStudy (anonymous):

can we combine the (x-3) and the (x) somehow with a property?